# Functional differential equations driven by càdlàg rough paths

### v1

@@ 2.2. Existence and uniqueness
$\|Y\|_{p,[\tau_{i},\tau_{i+1})}\lesssim_{p}(|y^{\prime}_{\tau_{i}}|+|F_{\tau_{i}}(Y)|+\|y^{\prime}\|_{p,[\tau_{i},\tau_{i+1})}+\|Y^{\prime}\|_{p,[\tau_{i},\tau_{i+1})})\|X\|_{p,[\tau_{i},\tau_{i+1}]}+\|R^{Y}\|_{\frac{p}{2},[\tau_{i},\tau_{i+1})}.$
Using (2.3) and (2.4), we can now conclude that there exist componentwise non-decreasing polynomials $Q_{p}^{(Y,J)},Q_{p}^{(Y^{\prime})}\colon[0,\infty)^{4}\to[0,\infty)$ such that
$|Y_{\tau_{i+1}-,\tau_{i+1}}(Y)|\leq Q_{p}^{(Y,J)}(|F_{\tau_{i}}(Y)|,\|y,y^{\prime}\|_{X,p,[\tau_{i},\tau_{i+1}]},C_{F},\|\mathbf{X}\|_{p,[\tau_{i},\tau_{i+1}]})$
as well as
(2.6) $\|Y^{\prime}\|_{p,[\tau_{i},\tau_{i+1}]}\leq Q_{p}^{(Y^{\prime})}(|F_{\tau_{i}}(Y)|,\|y,y^{\prime}\|_{X,p,[\tau_{i},\tau_{i+1}]},C_{F},\|\mathbf{X}\|_{p,[\tau_{i},\tau_{i+1}]}).$
So, finally, we need to control $|F_{\tau_{i}}(Y)|$ in terms of the information on $Y^{\prime}$, $R^{Y}$, etc. on the previous subinterval. It holds that
$|F_{\tau_{i}}(Y)|\leq|F_{\tau_{i-1}}(Y)|+C_{F}(1+\|Y\|_{p,[\tau_{i-1},\tau_{i}]})$
and
$\|Y\|_{p,[\tau_{i-1},\tau_{i}]}\lesssim_{p}(|y^{\prime}_{\tau_{i-1}}|+|F_{\tau_{i-1}}(Y)|+\|y^{\prime}\|_{p,[\tau_{i-1},\tau_{i}]}+\|Y^{\prime}\|_{p,[\tau_{i-1},\tau_{i}]})\|X\|_{p,[\tau_{i-1},\tau_{i}]}+\|R^{Y}\|_{\frac{p}{2},[\tau_{i-1},\tau_{i}]}.$
Thus, there exists a componentwise non-decreasing polynomial $Q_{p}^{(F(Y))}\colon[0,\infty)^{6}\to[0,\infty)$ such that
$|F_{\tau_{i}}(Y)|$
$\quad\leq Q_{p}^{(F(Y))}(|F_{\tau_{i-1}}(Y)|,\|y,y^{\prime}\|_{X,p,[\tau_{i-1},\tau_{i}]},C_{F},\|Y^{\prime}\|_{p,[\tau_{i-1},\tau_{i}]},\|R^{Y}\|_{\frac{p}{2},[\tau_{i-1},\tau_{i}]},\|\mathbf{X}\|_{p,[\tau_{i-1},\tau_{i}]}).$
Combining this with (2.5) and (2.6) in each iterative step, we obtain that
$|Y_{0}^{\prime}|+\|Y^{\prime}\|_{p}+\|R^{Y}\|_{\frac{p}{2}}\leq K_{p}(|F_{0}(y)|,\|y,y^{\prime}\|_{X,p},C_{F},\|\mathbf{X}\|_{p}),$
@@ 2.3. Continuity of the Itô–Lyons map
where the implicit multiplicative constant depends only on $p$, $C_{F}\vee C_{\tilde{F}}$, $C_{F,K,X,\tilde{X}}\vee C_{\tilde{F},K,X,\tilde{X}}$, $K$, $\|\mathbf{X}\|_{p}$ and $\|\tilde{\mathbf{X}}\|_{p}$.
@@
Step 1: Local estimate. Let $(Y,Y^{\prime})\in\mathcal{V}^{p}_{X}([0,t];\mathbb{R}^{k})$, $(\tilde{Y},\tilde{Y}^{\prime})\in\mathcal{V}^{p}_{\tilde{X}}([0,t];\mathbb{R}^{k})$ be local solutions to the RFDE (2.1), with data $((y,y^{\prime}),(F,F^{\prime}),\mathbf{X})$, $((\tilde{y},\tilde{y}^{\prime}),(\tilde{F},\tilde{F}^{\prime}),\tilde{\mathbf{X}})$, respectively, over a time interval $[0,t]$ for $t=t_{2}$ sufficiently small, see the proof of Theorem 2.3. Without loss of generality assume that $C_{\tilde{F}}\leq C_{F}$, $C_{\tilde{F},K,X,\tilde{X}}\leq C_{F,K,X,\tilde{X}}$. Assumption 2.1 gives that
@@
$\quad\leq C_{F,K,X,\tilde{X}}(|Y_{0}-\tilde{Y}_{0}|+\|Y-\tilde{Y}\|_{p,[0,t]})+C_{F-\tilde{F}}(1+\|\tilde{Y}\|_{p,[0,t]})$
@@
where the implicit multiplicative constant depends only on $p$, $\|\mathbf{X}\|_{p}$ and $\|\tilde{\mathbf{X}}\|_{p}$. Combining the results, we get that
@@
$\quad\leq\|y^{\prime}-\tilde{y}^{\prime}\|_{p,[0,t]}+\|F(Y)-\tilde{F}(\tilde{Y})\|_{p,[0,t]}+\|R^{y}-R^{\tilde{y}}\|_{\frac{p}{2},[0,t]}+\|R^{\int_{0}^{\cdot}F(Y)\mathrm{d}\mathbf{X}-\int_{0}^{\cdot}\tilde{F}(\tilde{Y})\mathrm{d}\tilde{\mathbf{X}}}\|_{\frac{p}{2},[0,t]}$
@@
for some constants $C_{1}>0$, which depends only on $p$, $C_{F,K,X,\tilde{X}}$, $\|\mathbf{X}\|_{p}$ and $\|\tilde{\mathbf{X}}\|_{p}$, and $C_{2}>1$, which depends additionally on $C_{F}$ and $K$. Hence, we can choose $t=t_{3}\leq t_{2}$ sufficiently small such that $C_{1}(\|\mathbf{X}\|_{p,[0,t]}\vee\|\tilde{\mathbf{X}}\|_{p,[0,t]})\leq\frac{1}{2}$, which implies that
@@
Step 2: Global estimate. We infer from Step 1 that there exists a constant $\gamma>0$, which depends only on $p$ and $C_{F,K,X,\tilde{X}}$, such that on any interval $[s,t]$ with $w(s,t)<\gamma$, for $w$ as introduced in the proof of Theorem 2.3, the local solutions satisfy an estimate of the form (2.3).
By [FZ18, Lemma 1.5], there exists a partition $\mathcal{P}=\{0=t_{0}<\dots<t_{N}=T\}$ of $[0,T]$, such that $w(t_{i},t_{i+1}-)<\gamma$ for every $i=0,1,\ldots,N-1$. Since $w$ is superadditive, we have that
$w(t_{0},t_{1}-)+w(t_{1}-,t_{1})+w(t_{1},t_{2}-)+\ldots+w(t_{N-1},t_{N}-)+w(t_{N}-,t_{N})\leq w(0,T).$
It is then straightforward to see that the partition $\mathcal{P}$ may be chosen such that the number of partition points $N$ may be bounded by a constant depending only on $\gamma$ and $w(0,T)$. Thus, we may combine the local estimates on each of the subintervals $[t_{i},t_{i+1})$, together with simple estimates on the jumps at the end-points of these subintervals, to obtain the global estimate. More precisely, we have the estimate
@@
for $i=0,\ldots,N-1$, where the implicit constant depends on $p$ and $C_{F,K,X,\tilde{X}}$, $C_{F}$, $K$, $\|\mathbf{X}\|_{p}$, $\|\tilde{\mathbf{X}}\|_{p}$, but not on the index $i$. So, it remains to bound
$|Y^{\prime}_{t_{i+1}-,t_{i+1}}-\tilde{Y}^{\prime}_{t_{i+1}-,t_{i+1}}|+|R^{Y}_{\tau_{i+1}-,\tau_{i+1}}-R^{\tilde{Y}}_{t_{i+1}-,t_{i+1}}|$
@@
Then, given the assumptions, and $(\int_{0}^{\cdot}F_{s}(Y)\,\mathrm{d}\mathbf{X}_{s})_{t-,t}=F_{t-}(Y)X_{t-,t}+F^{\prime}_{t-}(Y,Y^{\prime})\mathbb{X}_{t-,t}$, for $t\in(0,T]$, we have
@@
where the implicit multiplicative constant depends only on $p$, $C_{F}$, $C_{F,K,X,\tilde{X}}$, $K$, $\|\mathbf{X}\|_{p}$ and $\|\tilde{\mathbf{X}}\|_{p}$, and not on the index $i$. Here, we used Assumption (2.1) and the estimate (2.8) to derive that
@@
$\quad\leq|F^{\prime}_{t_{i+1}-}(Y,Y^{\prime})\mathbb{X}_{t_{i+1}-,t_{i+1}}-\tilde{F}^{\prime}_{t_{i+1}-}(\tilde{Y},\tilde{Y}^{\prime})\tilde{\mathbb{X}}_{t_{i+1}-,t_{i+1}}$
@@
$\quad\leq C_{F-\tilde{F}}(1+\|Y\|_{p,[t_{i},t_{i+1}]})$
@@
where the implicit multiplicative constant depends only on $p$, $C_{F}$, $C_{F,K,X,\tilde{X}}$, $K$ and $\|\mathbf{X}\|_{p}\allowbreak\vee\|\tilde{\mathbf{X}}\|_{p}$, and not on the index $i$.
@@
where the implicit multiplicative constant depends only on $p$, $C_{F}$, $C_{F,K,X,\tilde{X}}$, $K$, $\|X\|_{p}$ and $\|\tilde{X}\|_{p}$, and not on the index $i$; thus applying the estimate (2.3) for $i-1$ gives that
@@
$\qquad+\sum_{j=0}^{i-1}\Big{(}\|y^{\prime}-\tilde{y}^{\prime}\|_{p,[t_{j},t_{j+1}]}+\|R^{y}-R^{\tilde{y}}\|_{\frac{p}{2},[t_{j},t_{j+1}]}+C_{F-\tilde{F}}+\|\mathbf{X};\tilde{\mathbf{X}}\|_{p,[t_{j},t_{j+1}]}\Big{)}.$
Since $\|\cdot\|_{p,[0,T]}\leq N\sum_{i=0}^{N-1}\|\cdot\|_{p,[t_{i},t_{i+1}]}$ for any $p\geq 1$, this implies that
@@
Again, since $\|\cdot\|_{p,[0,T]}\leq N\sum_{i=0}^{N-1}\|\cdot\|_{p,[t_{i},t_{i+1}]}$ for any $p\geq 1$, the estimate finally follows. ∎
@@ 3.1. Classical RDEs
    where the implicit multiplicative constant depends only on $p$, $\|f\|_{C^{3}_{b}}\vee\|\tilde{f}\|_{C^{3}_{b}}$, $K$, $\|\mathbf{X}\|_{p}$ and $\|\tilde{\mathbf{X}}\|_{p}$.
@@
Growth conditions. It follows from the Lipschitz continuity of $f$ that
$\|F(Y)\|_{p,[s,t]}\leq\|f\|_{C^{2}_{b}}\|Y\|_{p,[s,t]}.$
Further, it holds that
@@
and
@@
Assumption 2.1 (i) therefore holds with some constant $C_{F}=\|f\|_{C^{2}_{b}}$ up to a multiplicative constant which depends only on $p$.
@@
(i) The existence and uniqueness of the solution follows immediately from Lemma 3.2 and Theorem 2.3. For the a priori estimate, note that $|F_{0}(y)|\leq\|f\|_{C^{2}_{b}}$, $\|y,y^{\prime}\|_{X,p}=0$ and $C_{F}\lesssim_{p}\|f\|_{C^{2}_{b}}$.
@@
satisfies Assumption 2.1 (i), since $C^{3}_{b}$ is a vector space. Thus we have, $C_{F-\tilde{F}}=\|f-\tilde{f}\|_{C^{3}_{b}}$ up to a multiplicative constant which depends only on $p$. ∎

### v2

@@ 2.2. Existence and uniqueness
(2.6) $\|Y^{\prime}\|_{p,[\tau_{i},\tau_{i+1}]}\leq Q_{p}^{(Y^{\prime})}(\|y,y^{\prime}\|_{X,p,[\tau_{i},\tau_{i+1}]},C_{F},\|\mathbf{X}\|_{p,[\tau_{i},\tau_{i+1}]}).$
Combining these estimates, we obtain that
$|Y_{0}^{\prime}|+\|Y^{\prime}\|_{p}+\|R^{Y}\|_{\frac{p}{2}}\leq K_{p}(\|y,y^{\prime}\|_{X,p},C_{F},\|\mathbf{X}\|_{p}),$
@@ 2.3. Continuity of the Itô–Lyons map
where the implicit multiplicative constant depends on $p$, $C_{F}\vee C_{\tilde{F}}$, $C_{F,K,X,\tilde{X}}\vee C_{\tilde{F},K,X,\tilde{X}}$, $K$, $\|\mathbf{X}\|_{p}$ and $\|\tilde{\mathbf{X}}\|_{p}$.
@@
Step 1: Local estimate. Let $(Y,Y^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$, $(\tilde{Y},\tilde{Y}^{\prime})\in\mathcal{V}^{p}_{\tilde{X}}([0,T];\mathbb{R}^{k})$ be the global solutions to the RFDE (2.2), with data $((y,y^{\prime}),(F,F^{\prime}),\mathbf{X})$, $((\tilde{y},\tilde{y}^{\prime}),(\tilde{F},\tilde{F}^{\prime}),\tilde{\mathbf{X}})$, respectively, see Theorem 2.4. Let $t\in(0,T]$. Without loss of generality assume that $C_{\tilde{F}}\leq C_{F}$, $C_{\tilde{F},K,X,\tilde{X}}\leq C_{F,K,X,\tilde{X}}$. Assumption 2.1 gives that
@@
$\quad\leq C_{F,K,X,\tilde{X}}(|Y_{0}-\tilde{Y}_{0}|+\|Y-\tilde{Y}\|_{p,[0,t]})$
$\qquad+C_{F-\tilde{F}}(1+(|\tilde{Y}^{\prime}_{0}|+\|\tilde{Y}^{\prime}\|_{p,[0,t]})\|X\|_{p,[0,t]}+\|R^{\tilde{Y}}\|_{\frac{p}{2},[0,t]})$
@@
where the implicit multiplicative constant depends on $p$, $\|\mathbf{X}\|_{p}$ and $\|\tilde{\mathbf{X}}\|_{p}$. Combining the results, we get that
@@
$\quad\leq\|y^{\prime}-\tilde{y}^{\prime}\|_{p,[0,t]}+\|F(Y)-\tilde{F}(\tilde{Y})\|_{p,[0,t]}$
$\qquad+\|R^{y}-R^{\tilde{y}}\|_{\frac{p}{2},[0,t]}+\|R^{\int_{0}^{\cdot}F(Y)\mathrm{d}\mathbf{X}-\int_{0}^{\cdot}\tilde{F}(\tilde{Y})\mathrm{d}\tilde{\mathbf{X}}}\|_{\frac{p}{2},[0,t]}$
@@
for some constants $C_{1}>0$, which depends on $p$, $C_{F,K,X,\tilde{X}}$, $\|\mathbf{X}\|_{p}$ and $\|\tilde{\mathbf{X}}\|_{p}$, and $C_{2}>1$, which depends additionally on $C_{F}$ and $K$. Hence, we can choose $t$ sufficiently small such that $C_{1}(\|\mathbf{X}\|_{p,[0,t]}\vee\|\tilde{\mathbf{X}}\|_{p,[0,t]})\leq\frac{1}{2}$, which implies that
@@
Step 2: Global estimate. Recall the right-continuous control function $w\colon\Delta_{T}\to[0,\infty)$ given by
$w(s,t):=\|X\|_{p,[s,t]}^{p}+\|\mathbb{X}\|_{\frac{p}{2},[s,t]}^{\frac{p}{2}},\quad(s,t)\in\Delta_{T},$
as introduced in the proof of Theorem 2.4, and let $\tilde{w}(s,t):=\|\tilde{X}\|_{p,[s,t]}^{p}+\|\tilde{\mathbb{X}}\|_{\frac{p}{2},[s,t]}^{\frac{p}{2}}$, $(s,t)\in\Delta_{T}$. We infer from Step 1 that there exists a constant $\gamma>0$, which depends on $p$, $C_{F,K,X,\tilde{X}}$, $\|\mathbf{X}\|_{p}$ and $\|\tilde{\mathbf{X}}\|_{p}$, such that on any interval $[s,t]$ with $w(s,t)\vee\tilde{w}(s,t)\leq\gamma$ the local solutions satisfy an estimate of the form (2.3).
Let $c(s,t):=w(s,t)+\tilde{w}(s,t)$, $(s,t)\in\Delta_{T}$. Since $c$ is right-continuous, there exists a partition $\mathcal{P}=\{0=t_{0}<\dots<t_{N}=T\}$ of $[0,T]$, such that
$c(t_{i},t_{i+1}-)=\gamma,\quad\text{or}\quad c(t_{i},t_{i+1}-)<\gamma\quad\text{and}\quad c(t_{i},t_{i+1}-)+c(t_{i+1}-,t_{i+1})\geq\gamma,$
for every $i=0,1,\ldots,N-1$. Since $w$ and $\tilde{w}$ and, thus, $c$ is superadditive, we have that
$N\gamma\leq\sum_{i=0}^{N-1}c(t_{i},t_{i+1}-)+c(t_{i+1}-,t_{i+1})\leq c(0,T).$
Therefore, the number of partition points $N$ may be bounded by a constant depending only on $\gamma$, $w(0,T)$ and $\tilde{w}(0,T)$. Thus, in this step, we may combine the local estimates on each of the subintervals $[t_{i},t_{i+1})$, together with simple estimates on the jumps at the end-points of these subintervals, which we aim to derive, to obtain the global estimate. More precisely, by Step 1, we have the local estimate
@@
for $i=0,\ldots,N-1$, where the implicit multiplicative constant depends on $p$ and $C_{F,K,X,\tilde{X}}$, $C_{F}$, $K$, $\|\mathbf{X}\|_{p}$, $\|\tilde{\mathbf{X}}\|_{p}$, but not on the index $i$. So, it remains to bound
$|Y^{\prime}_{t_{i+1}-,t_{i+1}}-\tilde{Y}^{\prime}_{t_{i+1}-,t_{i+1}}|+|R^{Y}_{t_{i+1}-,t_{i+1}}-R^{\tilde{Y}}_{t_{i+1}-,t_{i+1}}|$
@@
We note that $(\int_{0}^{\cdot}F_{s}(Y)\,\mathrm{d}\mathbf{X}_{s})_{t-,t}=F_{t-}(Y)X_{t-,t}+F^{\prime}_{t-}(Y,Y^{\prime})\mathbb{X}_{t-,t}$, that is, with $Y_{t-,t}=y_{t-,t}+(\int_{0}^{\cdot}F_{s}(Y)\,\mathrm{d}\mathbf{X}_{s})_{t-,t}$, it follows that $Y^{\prime}_{t-,t}=y^{\prime}_{t-,t}+F_{t-,t}(Y)$ and $R^{Y}_{t-,t}=R^{y}_{t-,t}+F^{\prime}_{t-}(Y,Y^{\prime})\mathbb{X}_{t-,t}$, for $t\in(0,T]$. Given the assumptions, we then have
@@
where the implicit multiplicative constant depends on $p$, $C_{F}$, $C_{F,K,X,\tilde{X}}$, $K$, $\|\mathbf{X}\|_{p}$ and $\|\tilde{\mathbf{X}}\|_{p}$, and not on the index $i$. Here, we used Assumption (2.1) and the estimate (2.8) to derive that
@@
$\quad\leq|F^{\prime}_{t_{i+1}-}(Y,Y^{\prime})\mathbb{X}_{t_{i+1}-,t_{i+1}}-\tilde{F}^{\prime}_{t_{i+1}-}(\tilde{Y},\tilde{Y}^{\prime})\tilde{\mathbb{X}}_{t_{i+1}-,t_{i+1}}|$
@@
$\quad\leq C_{F-\tilde{F}}(1+(|Y^{\prime}_{t_{i}}|+\|Y^{\prime}\|_{p,[t_{i},t_{i+1}]})\|X\|_{p,[t_{i},t_{i+1}]}+\|R^{Y}\|_{\frac{p}{2},[t_{i},t_{i+1}]})$
@@
where the implicit multiplicative constant depends on $p$, $C_{F}$, $C_{F,K,X,\tilde{X}}$, $K$ and $\|\mathbf{X}\|_{p}\allowbreak\vee\|\tilde{\mathbf{X}}\|_{p}$, and not on the index $i$.
@@
where the implicit multiplicative constant depends on $p$, $C_{F}$, $C_{F,K,X,\tilde{X}}$, $K$, $\|X\|_{p}$ and $\|\tilde{X}\|_{p}$, and not on the index $i$; thus applying the estimate (2.3) for $i-1$ gives that
@@
$\qquad+\sum_{j=0}^{i-1}\Big{(}\|y^{\prime}-\tilde{y}^{\prime}\|_{p,[t_{j},t_{j+1}]}+\|R^{y}-R^{\tilde{y}}\|_{\frac{p}{2},[t_{j},t_{j+1}]}+C_{F-\tilde{F}}+\|\mathbf{X};\tilde{\mathbf{X}}\|_{p,[t_{j},t_{j+1}]}\Big{)},$
that is,
$(|Y_{t_{i}}-\tilde{Y}_{t_{i}}|+|F_{t_{i}}(Y)-\tilde{F}_{t_{i}}(\tilde{Y})|+|y^{\prime}_{t_{i}}-\tilde{y}^{\prime}_{t_{i}}|)^{p}$
$\quad\lesssim(|y_{0}-\tilde{y}_{0}|+|F_{0}(y)-\tilde{F}_{0}(\tilde{y})|+|y^{\prime}_{0}-\tilde{y}^{\prime}_{0}|+NC_{F-\tilde{F}})^{p}$
$\qquad+\sum_{j=0}^{i-1}\Big{(}\|y^{\prime}-\tilde{y}^{\prime}\|_{p,[t_{j},t_{j+1}]}^{p}+\|R^{y}-R^{\tilde{y}}\|_{\frac{p}{2},[t_{j},t_{j+1}]}^{p}+\|\mathbf{X};\tilde{\mathbf{X}}\|_{p,[t_{j},t_{j+1}]}^{p}\Big{)}.$
This implies that
@@
Since $\|\cdot\|_{p,[0,T]}\leq N\sum_{i=0}^{N-1}\|\cdot\|_{p,[t_{i},t_{i+1}]}$ for any $p\geq 1$, see e.g. [All21, Lemma A.1], the estimate finally follows. ∎
@@ 3.1. Classical RDEs
    where the implicit multiplicative constant depends on $p$, $\|f\|_{C^{3}_{b}}\vee\|\tilde{f}\|_{C^{3}_{b}}$, $K$, $\|\mathbf{X}\|_{p}$ and $\|\tilde{\mathbf{X}}\|_{p}$.
@@
Growth conditions. It is clear that $|F_{t}(Y)|\leq\|f\|_{C^{2}_{b}}$, and it follows from the Lipschitz continuity of $f$ that
$|F_{t-,t}(Y)|\leq\|F(Y)\|_{p,[s,t]}\leq\|f\|_{C^{2}_{b}}\|Y\|_{p,[s,t]}.$
We now note that $\|Y\|_{p,[s,t]}\leq\|Y\|_{p,[s,t)}+|Y_{t-,t}|$ as well as $\|Y\|_{p,[s,t]}\leq C_{p}(1+(|Y^{\prime}_{s}|+\|Y^{\prime}\|_{p,[s,t]})\cdot\|X\|_{p,[s,t]}+\|R^{Y}\|_{\frac{p}{2},[s,t]})$. Further, it holds that
@@
and by Taylor’s expansion,
$R^{F(Y)}_{s,t}$ $=R^{f(Y)}_{s,t}=f(Y_{t})-f(Y_{s})-\mathrm{D}f(Y_{s})Y_{s,t}+\mathrm{D}f(Y_{s})R^{Y}_{s,t}$
$=\frac{1}{2}\mathrm{D}^{2}f(Y_{s}+\lambda Y_{s,t})Y^{2}_{s,t}+\mathrm{D}f(Y_{s})R^{Y}_{s,t},$
with $\lambda\in[0,1]$, which implies that
@@
Assumption 2.1 (i) therefore holds with some constant $C_{F}=\|f\|_{C^{2}_{b}}$ up to a multiplicative constant which depends on $p$.
@@
(i) The existence and uniqueness of the solution follows immediately from Lemma 3.2 and Theorem 2.4. For the a priori estimate, note that $\|y,y^{\prime}\|_{X,p}=0$ and $C_{F}\lesssim_{p}\|f\|_{C^{2}_{b}}$.
@@
satisfies Assumption 2.1 (i), since $C^{3}_{b}$ is a vector space. Thus we have, $C_{F-\tilde{F}}\lesssim_{p}\|f-\tilde{f}\|_{C^{2}_{b}}$. ∎